A box contains 10 green balls and 8 red balls. Six balls are selected at random. Find the probability that:

a) all 6 balls are green.

b) exactly 4 out of the 6 balls are red.

c) at least one of the 6 balls is red.

d) 3 or 4 of the 6 balls are red.


1
Expert's answer
2021-07-25T17:09:29-0400
"10+8=18(balls)"

a)

"P(6\ green)=\dfrac{\dbinom{10}{6}\dbinom{8}{0}}{\dbinom{18}{6}}=\dfrac{210(1)}{18564}=\dfrac{5}{442}"

"\approx0.011312"

(b)


"P(4\ red)=\dfrac{\dbinom{10}{2}\dbinom{8}{4}}{\dbinom{18}{6}}=\dfrac{45(70)}{18564}=\dfrac{75}{442}"


"\approx0.169683"



(c)


"P(at\ least\ 1\ red)=1-P(6\ green)"

"=1-\dfrac{5}{442}=\dfrac{437}{442}"

"\approx0.988682"

(d)


"P(3\ red\ or\ 4\ red)=P(3\ red)+P(4\ red)"


"=\dfrac{\dbinom{10}{3}\dbinom{8}{3}}{\dbinom{18}{6}}+\dfrac{\dbinom{10}{2}\dbinom{8}{4}}{\dbinom{18}{6}}"

"=\dfrac{120(56)}{18564}+\dfrac{45(70)}{18564}=\dfrac{235}{442}"

"\approx0.525727"